Ask
your own question, for FREE!
Probability
8 Online
An urn contains 6 red balls and 3 blue balls. One ball is selected at random and is replaces by a ball of the= other color. A Second ball is then chosen. What is the conditional probability that the first all selected is red, given that the second ball was red?
Still Need Help?
Join the QuestionCove community and study together with friends!
R_1 - First ball Red, R_2-second ball is red P(R_1|R_2)=P(R_1 AND R_2)/P(R_2) P(R_1 AND R_2) = 6/9 * 5/9 = 30/81=10/27 P(R_2) = 6/9 * 5/9 + 3/9 * 6/9 = 16/27 P(R_1|R_2)=(10/27)/(16/27) = 5/8. I got it wrong. Why?
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
1 day ago
1 Reply
0 Medals
5 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
3 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal